Y PC PB YD TEMP PRP
39.7 42.3 143.8 50.1 -16 107.8
38.69 49.4 152.2 54.98 -4 134.6
42.02 45.5 145.7 59.72 -24 134
42.71 45.3 145.9 65.17 16 125.4
44.75 49.3 178.8 72.24 5 143.6
48.35 50 222.4 79.67 13 152.5
48.47 53.5 233.6 88.22 21 147.5
50.37 53.8 234.7 97.65 49 161.2
51.52 51.5 238.4 104.26 4 185.6
52.55 56 234.1 111.31 35 179.7
54.61 61.5 235.5 123.19 11 171.4
56.42 56.2 228.6 130.37 4 170.8
57.7 63.1 226.8 136.49 18 188.8
61.94 53.1 238.4 142.41 35 199.4
63.8 62.1 250.3 152.97 46 194
66.88 64.2 265.7 162.57 32 193.5
70.34 60.5 281 171.31 64 224.9
73.26 57.7 288.3 176.09 52 224.2
76.39 59 284.6 184.94 18 209.5
78.27 27.1 293.4 188.72 27 209.1
79.65 26.2 282.9 195.55 48 209.5
79.27 26.9 284.3 202.87 71 206.1
80.61 28 280.2 210.91 36 233.7
83.1 33.2 279.5 219.4 60 245
83.76 33.4 277.1 231.61 89 242.7
88.98 39.5 287.8 239.68 60 241.4
90.08 43 306.4 254.69 62 258.2
89.71 43.4 337.7 262.24 74 269.4
94.37 43.9 331.5 271.45 85 265.8

Download the data set CHICK.DTA   

  1. Estimate demand for per capita chicken consumption Y (in pounds) as a linear function of its price PC (in cents), price of its substitute beef PB, and the consumer disposable income YD. Assess the results.
  2. Re-estimate the equation without the variable PB. Comment on the results by comparison with those in a. above.
  3. Now re-estimate the equation in a. by adding a new variable for temperature TEMP. Comment on the results by comparison with those in a. above.
Answer #1

The required estimations can be done by simple R-commands. The command and output is shown as below. To input the data in R, use the following code.

-----------------------------------------------

> library(readr)
> dat <- read_delim("dat", "\t", escape_double = FALSE, trim_ws = TRUE)

-----------------------------------------------

(a) The regression output would be as below.

> summary(Im(Y - PC + PB + YD, data = dat)) Call: lm(formula = Y - PC + PB + YD, data = dat), Residuals: Min 10 Median 30 Ma

Except the variable PB, all the coefficients are individually statistically significant at 1% (highly significant). However, if the significance level is shifted to 10%, then all the coefficients are significant. What can be stated is that the variable PB is less significant than other variables.

The R-squared is high, and F-statistic (for the test of overall significance) is also highly significant (p-value much less than 1%). This means that the model explains the dependent variable very well.

(b) The regression output would be as below.

> summary(lm(Y - PC + YD, data = dat)) Call: ilm(formula = Y - PC + YD, data = dat), Residuals: Min 1Q Median -3.2288 -0.9927

In this case, all the coefficients are individually statistically significant at 5%. But at 1%, PC coefficient is not significant (since p-value greater than 1%).

As can be suspected, the R-square and adjusted R-square both decreased as we have removed a variable. But the F-statistic is certainly significant again.

(c) The regression output would be as below.

> summary(lm(Y - PC + PB + YD + TEMP, data = dat)) Call: lm( formula = Y - PC + PB + YD + TEMP, data = dat), Residuals: Min 1

The coefficients of variable PB and TEMP are not statistically significant at 1% or even at 5%. Also, TEMP is not at all significant at 10% too. Rest of the variables are highly significant as before.

The R-square marginally increased, but the adjusted R-square is same as before. The F-statistic is also significant, as before.

All that can be said is that, the TEMP and PB variable must be tested via restricted and unrestricted regression, to be included in the model.